Click Submit to complete this dsessment. Question 10 Maximising the value of the firm can also be described as: maximising the value of the companys ordinary shares. maximising the accounting profit of the firm. minimising the interest rate charged by creditors of the firm. maximising sales revenue or turnover of the firm. A Click Submit to complete this assessment.

Remaining Time: 59 minutes, 33 seconds. v Question Completion Status Δ Moving to the next question prevents changes to this answer. Question > Question1 1 points S You have just won a lottery, which offers a choice in how you may receive your prize. If the interest rate is 10% per compounded annually, which of the following prizes has the highest present value? O $2 000 000 immediately An ordinary annuity of $300 000 per year for each of the next 10 years (starting one year from now) o $3 000 000 at the end of five years A perpetuity of $250 000 per year starting one year from now Moving to the next question prevents changes to this answer. Ques
Answer #1

10:

Maximizing the value of the company's ordinary shares

(The value of firm is described as value of the company. The accounting profit/sales of the firm does not determine the value. Interest charged by creditors is not a determinant of value of firm)

1:

A perpetuity of $250,000 starting 1 year from now

PV of first option = $2000,000

PV of option 2 = A*(1-1/(1+r)^n)/r

= 300000*(1-1/1.1^10)/0.1

=1,843,370

PV of option 3= Fv/(1+r)^n

= 3000000/1.1^5

=1862764

PV of option 4= (P/r)/ (1+r)^1

= (250000/0.1)/1.1

=2,272,727

Option 4 has the highest PV